You pick a card from a deck. If it is a face card, you will win 500. If you get an ace, you will win 1000. If the card you picked is red you get 100. For any other card, you will win nothing. Find the expected value that you can possibly win

1
Expert's answer
2022-05-30T05:29:17-0400

There are 52 cards in the deck.

There are 12 face cards.

There are 4 aces.

There are 26 red cards.


"E(X)=\dfrac{12}{52}(500)+\dfrac{4}{52}(1000)+\dfrac{26}{52}(100)=242.31"